# Reducing Spatial Discretization Error on Coarse CFD Simulations Using an OpenFOAM-Embedded Deep Learning Framework

### v2

@@ 5.3 Generalization
Once we checked the model’s good behavior under the training data, we then evaluate its generalization outside the training distribution. Theoretically, since the model exploits the local features of the physics, it can be applied to any flow with characteristics similar to those seen during training. For this purpose, we prepare two experiments: (1) extending the simulation time of the original case until $t=25$ s, and (2) applying the model to a case with two tandem columns with an extended domain.
@@
From Figure 15(a) and Figure 15(b), we observe that both components of the velocity follow a similar trend, reducing the DC model error by about 50%. In the case of the pressure, both the DC and the baseline solver get a similar error, continuing the behavior of the training data. For the turbulent variables, from Figure 15(d) and Figure 15(e), we conclude that the DC model reduces the error compared to the baseline solver, although in a minor grade than with the training data.
@@
Although the tandem experiment´s flow regime is the same as in training data, the second column introduces new fluid behavior, becoming a more challenging experiment than just expanding the domain. Despite this difficulty, Figures 16 (a) and (b) show that the Deep-Convection model reduces the error in both components of the velocity by around 50% compared with the baseline. The reduction in the error is more evident after $t=14$ s, where the wake is fully developed in both columns, showing the same trend as with the training data. Thus, we expect a higher reduction in the error before $t=20$ s, both for the accumulation of the error in the baseline solver and for the full development of the flow behind both columns. Figure 16 (c) shows agreement in the pressure with the previous results, with no significant variation compared with the baseline solver. The turbulent quantities follow a similar reduction than in the time-extended experiment, with a higher improvement expected in the following timesteps due to the total formation of the wakes. For a deeper exploration of the results, we include the evolution of some variables along the longitudinal line $y=3D/4$ (taking $y=0$ the lower wall at the bottom of the domain) at $t=20$ s in Figure 17. In particular, we observe a better agreement of the velocity generated by the DC model with the ground truth and a more precise concordance of the DC model to the complex evolution of the turbulence viscosity $\nu_{t}$, defined as the ratio between $k$ and $\omega$.
@@
The presented results show that the model accomplishes its objective of reducing the spatial discretization error in coarse meshes. Our objective was to apply the model to the fully turbulent regime, so while maintaining the Reynolds number of the simulation close to the Reynolds number of the training data, the model is expected to generalize well. In the case of extending the model to different Reynolds numbers, it would be necessary to include some knowledge of the turbulence in the input of the neural network (e.g., turbulence viscosity of the Reynolds number). In the case of long-term simulations, the model is expected to produce degraded results due to the common problem of neural networks predicting time series. A solution could be to monitor the residual of the NSE and switch to the baseline solver when the residual is over a certain threshold, as done in [17].
6 Conclusions
@@
We have developed and validated an OpenFOAM-embedded deep learning framework for incorporating fine-scale effects and flow structures into coarse mesh numerical simulations, minimizing the spatial discretization error at a reduced computational cost. We employed a feed-forward neural network to replace the traditional differencing scheme for the convective term. The model learns from simulations generated with a fine mesh and modifies the coarse solver to generate results close to the trained data. The framework is an end-to-end differentiable model, so we employed a discrete adjoint version of the CFD code to differentiate the physics automatically. We also developed a fast communication method between the programming languages of the programs (Python and C++). We trained the model with fine data for a flow past a square cylinder problem at a Reynolds number $Re=5\cdot 10^{5}$, obtaining 60 times faster results with a 25% error for both velocity components. We assessed the model’s generalization outside the training distribution in a time-extended and a two-columns-in-tandem simulation. The model’s response was similar in both cases, generating an error reduction of 50% for both components of the velocity, a slight reduction in the turbulent variables, and a similar error in the pressure, compared with the baseline model. Hence, our model brings a new tool for generating precise results at a significantly reduced computational cost with application to a broad range of fluid dynamics problems.
@@ Acknowledgement
The authors have received funding from the Spanish Ministry of Science and Innovation projects with references TED2021-132783B-I00 funded by MCIN/ AEI /10.13039/501100011033 and the European Union NextGeneration EU/ PRTR, PID2019-108111RB-I00 funded by MCIN/ AEI /10.13039/501100011033, the “BCAM Severo Ochoa” accreditation of excellence CEX2021-001142-S / MICIN / AEI / 10.13039/501100011033; the Spanish Ministry of Economic and Digital Transformation with Misiones Project IA4TES (MIA.2021.M04.008 / NextGenerationEU PRTR); and the Basque Government through the BERC 2022-2025 program, the Elkartek projects BEREZ-IA (KK-2023/00012) and EKIOCEAN (KK-2023/00097), and the Consolidated Research Group MATHMODE (IT1456-22) given by the Department of Education.
@@ Appendix A Construction of training data from a CFD simulation
After performing a CFD simulation with the features described in sections 2 and 4.4, we obtain $n$ snapshots, one for each timestep $t_{i}\in[t_{0},t_{1},...,t_{n}]$. We up-sample this fine data to the desired coarse resolution using the projection operator $\mathcal{P}_{f}^{c}$. Thus, we generate a projected version of the fine data that gathers each involved variable $(\bar{\textbf{u}},\bar{p},k,\omega)$ at time $t_{i}$, forming the training snapshots $d_{i}\in[d_{0},d_{1}...,d_{n}]$.

### v3

@@ 5.3 Generalization
Once we verified the model’s good behavior under the training data, we then evaluate its generalization outside the training distribution. Theoretically, since the model exploits the local features of the physics, it can be applied to any flow with characteristics similar to those seen during training. For this purpose, we prepare two experiments: (1) extending the simulation time of the original case until $t=25$ s, and (2) applying the model to a case with two tandem columns with an extended domain.
@@
Figures 16(a) and 16(b) show that both components of the velocity follow a similar trend, reducing the DC model error by about 50%. In the case of the pressure, both the DC and the baseline solver get a similar error, continuing the behavior of the training data. For the turbulent variables, Figures 16(d) and 16(e) show that the DC model reduces the error compared to the baseline solver, although in a minor grade than with the training data.
@@
Although the tandem experiment´s flow regime is the same as in training data, the second column introduces new fluid behavior, becoming a more challenging experiment than just expanding the domain. Despite this difficulty, Figures 17(a) and 17(b) show that the Deep-Convection model reduces the error in both components of the velocity by around 50% compared with the baseline. The reduction in the error is more evident after $t=14$ s, where the wake is fully developed in both columns, showing the same trend as with the training data. Thus, we expect a higher reduction in the error before $t=20$ s, both for the accumulation of the error in the baseline solver and for the full development of the flow behind the columns. Figure 17(c) shows agreement in the pressure with the previous results, with no significant variation compared with the baseline solver. The turbulent quantities follow a similar reduction than in the time-extended experiment, with a higher improvement expected in the following timesteps due to the total formation of the wakes. For a deeper exploration of the results, we include the evolution of some variables along the longitudinal line $y=3D/4$ (taking $y=0$ the lower wall at the bottom of the domain) at $t=20$ s in Figure 18. In particular, we observe a better agreement of the velocity generated by the DC model with the ground truth and a more precise concordance of the DC model to the complex evolution of the turbulence viscosity $\nu_{t}$, defined as the ratio between $k$ and $\omega$.
@@
These results show that the model accomplishes its objective of reducing the spatial discretization error in coarse meshes. Our model is unaware of the flow regime for two reasons: it receives the inputs in a normalized way and it does not see any information about the Reynolds number. Hence, it is unable to generalize to other Reynodls numbers different from the one it was trained for. This limitation is inherent to our model since it was not conceived for that purpose. Our objective was to apply our method to the fully turbulent regime, so while maintaining the Reynolds number of the simulation close to the Reynolds number of the training data, the model is expected to generalize well. In the case of extending its application to different Reynolds numbers, it would be necessary to include some knowledge of the grade of turbulence in the input of the neural network (e.g., turbulent viscosity of the Reynolds number). In the case of long-term simulations, the model is expected to produce degraded results due to the common problem of neural networks predicting time series. We extended the simulation of the two-columns-in-tandem in time to illustrate this behavior (see Figure 19). A solution could be to monitor the residual of the NSE and switch to the baseline solver when the residual is over a certain threshold, as done in [17].
Figure 19: Comparison of results for the two-columns-in-tandem experiment extending the simulation time until failure. (a) Error on the x-component of the velocity. (b) Error on the y-component of the velocity. (c) Error on the pressure. (d) Error on the $k$ variable. (e) Error on the $\omega$ variable.
6 Conclusions and Future Work
@@
We have developed and validated an OpenFOAM-embedded deep learning framework for incorporating fine-scale effects and flow structures into coarse mesh numerical simulations, minimizing the spatial discretization error at a reduced computational cost. We employed a feed-forward neural network to replace the traditional differencing scheme for the convective term. The model learns from simulations generated with a fine mesh and modifies the coarse solver to generate results close to the trained data. The framework is an end-to-end differentiable model, so we employed a discrete adjoint version of the CFD code to differentiate the physics automatically. We also developed a fast communication method between the programming languages of the programs (Python and C++). We trained the model with fine data projected to the coarse dimension for a flow past a square cylinder problem at a Reynolds number $Re=5\cdot 10^{5}$, obtaining 60 times faster results with a 25% error for both velocity components. We assessed the model’s generalization outside the training distribution in a time-extended and a two-columns-in-tandem simulation. The model’s response was similar in both cases, generating an error reduction of 50% for both components of the velocity, a slight reduction in the turbulent variables, and a similar error in the pressure, compared with the baseline model. Hence, our model brings a new tool for generating precise results at a significantly reduced computational cost with application to a broad range of fluid dynamics problems.
As possible future lines of research, we highlight the inclusion of a second neural network to correct the pressures and the extension of the method to LES simulations, with the objective of improving the drag prediction. Another research line would be to extend the technique to unstructured meshes, considering new architectures like graph neural networks. Further research on the influence of the down-sampling operator would be highly desirable to understand the method’s capabilities. Finally, following the recent developments in uncertainty quantification in fluid dynamics [61], we find it promising to combine the presented methodology with machine-learning-based uncertainty quantification for both aleatoric (data) [62] and epistemic (models) [63] uncertainties.
@@ Acknowledgement
The authors have received funding from the Spanish Ministry of Science and Innovation projects with references TED2021-132783B-I00 funded by MCIN/ AEI /10.13039/5011000110
33 and the European Union NextGeneration EU/ PRTR, PID2019-108111RB-I00 funded by MCIN/ AEI /10.13039/501100011033, PID2023-146678OB-I00 funded by MICIU/AEI/10.130
39/501100011033 and by the European Union NextGenerationEU/ PRTR, the “BCAM Severo Ochoa” accreditation of excellence CEX2021-001142-S / MICIN / AEI / 10.13039/5011
00011033, PRE2020-093093 funded by MICIU/AEI/10.13039/501100011033 and EI ESF “ESF Investing in your future”; the Spanish Ministry of Economic and Digital Transformation with Misiones Project IA4TES (MIA.2021.M04.008 / NextGenerationEU PRTR); and the Basque Government through the BERC 2022-2025 program, the Elkartek projects BEREZ-IA (KK-2023/00012) and EKIOCEAN (KK-2023/00097), and the Consolidated Research Group MATHMODE (IT1456-22) given by the Department of Education.
@@ Appendix A Construction of training data from a CFD simulation
After performing a CFD simulation with the features described in sections 2 and 4.4, we obtain $n$ snapshots, one for each timestep $t_{i}\in[t_{0},t_{1},...,t_{n}]$. We down-sample this fine data to the desired coarse resolution using the projection operator $\mathcal{P}_{f}^{c}$. Thus, we generate a projected version of the fine data that gathers each involved variable $(\bar{\textbf{u}},\bar{p},k,\omega)$ at time $t_{i}$, forming the training snapshots $d_{i}\in[d_{0},d_{1}...,d_{n}]$.